40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14503     Entry time: Sun Mar 31 15:05:53 2019     In reply to: 14502     Reply to this: 14510   14516   14643
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fiber beam-splitters not PM 

I looked into this a little more today.

  1. Looking at the beat signal between the PSL and EX beams from the NF1611 on a scope (50-ohm input), the signal Vpp was ~200 mV.
  2. In the time that I was poking about, the level dropped to ~150mVpp. seemed suspicious.
  3. Thinking that this has to be related to the polarization mismatch between the interfering beams, I moved the input fibers (blue in Attachment #1) around, and saw the signal amplitude went up to 300mVpp, supporting my initial hypothesis.
  4. The question remains as to where the bulk of the polarization drift is happening. I had spent some effort making sure the input coupled beam to the fiber was well-aligned to one of the special axes of the fiber, and I don't think this will have changed since (i.e. the rotational orientation of the fiber axes relative to the input beam was fixed, since we are using the K6XS mounts with a locking screw for the input couplers). So I flexed the patch cables of the fiber beam splitters inside the BeatMouth, and saw the signal go as high as 700mVpp (the expected level given the values reported by the DC monitor).

This is a problem - such large shifts in the signal level means we have to leave sufficient headroom in the choice of RF amplifier gain to prevent saturation, whereas we want to boost the signal as much as possible. Moreover, this kind of operation of tweaking the fiber seating to increase the RF signal level is not repeatable/reliable. Options as I see it:

  1. Get a fiber BS that is capable of maintaining the beam polarization all the way through to the beat photodiode. I've asked AFW technologies (the company that made our existing fiber BS parts) if they supply such a device, and Andrew is looking into a similar component from Thorlabs.
    • These parts could be costly.
  2. Mix the beams in free space. We have the beam coming from EX to the PSL table, so once we mix the two beams, we can use either a fiber or free-space PD to read out the beatnote. 
    • This approach means we lose some of the advantages of the fiber based setup (e.g. frequent alignment of the free-space MM of the two interfering beams may be required).
    • Potentially increases sensitivity to jitter noise at the free-space/fiber coupling points
Quote:
    • An initial RF beat power level measurement yielded -5dBm, which is inconsistent with the DC monitor voltages, but I'm not sure what frequency the beat was at, will make a more careful measurement with a scope or the network analyzer.
Attachment 1: IMG_7384.JPG  1.248 MB  Uploaded Sun Mar 31 16:21:24 2019  | Hide | Hide all
IMG_7384.JPG
ELOG V3.1.3-